# Rectification Reimagined: A Unified Mamba Model for Image Correction and Rectangling with Prompts

### v1

@@ Appendix A Supplementary Materials for UniRect
In this section, we discuss the various tasks under our proposed general distortion model. The general distortion model incorporates the Kannala-Brandt model (kannala2006generic), typically used to approximate wide-angle lenses, and the Brown-Conrady model (weng1992camera), employed for radial distortion modeling. Consequently, we can employ this proposed distortion model to simulate real image distortion in real-world smartphone camera systems. As shown in Eq. (3), This model can be simplified as
@@
Optical flow is a concept related to the analysis of the movement of objects in a sequence of images. We utilize this tool to explore the distortion process for all pixels and formulate each distortion for every task. This is a qualitative tool. The generation of optical flow via different methods is presented in Fig. 11. Taking the distortion $\mathcal{T}_{2}^{-1}$ as an example, we present the results of various optical flow estimation methods, such as FlowNet (dosovitskiy2015flownet), FlowNet2 (ilg2017flownet), PWC-Net (sun2018pwc), and RAFT (teed2020raft). It can be observed from the figure that the outcomes of different optical flow estimation methods are similar without altering the mathematical form of distortion derived in this paper.
@@
Dataset Construction. For portrait correction $\mathcal{T}_{1}$, the data was collected using five different ultra wide-angle smartphone cameras. Over ten people were photographed in several scenes (tan2021practical). The wide - angle portraits were manually corrected into distortion-free images. For the rectified wide-angle image rectangling dataset (liao2023recrecnet), wide-angle images were first synthesized, and the ground-truth was obtained manually. In the stitched image rectangling dataset (nie2022deep) utilized for $\mathcal{T}_{3}$, the majority of the inputs are real stitching results, while the ground - truth is synthesized manually. For rotation correction $\mathcal{T}_{4}$ (nie2023deep), the images are sourced from ImageNet (deng2009imagenet) and processed manually.
@@
Zhu et al. (zhu2022semi) $\mathcal{T}_{1}$ 8.79M - -
RecRecNet (liao2023recrecnet) $\mathcal{T}_{2}$ 62.70M - -
Nie et al. (nie2022deep) $\mathcal{T}_{3}$ 52.14M - 20
RecDiffusion (zhou2024recdiffusion) $\mathcal{T}_{3}$ 800M >1T -
DRC (nie2023deep) $\mathcal{T}_{4}$ - - 5
@@
Here, we introduce the details of two metrics (i.e. LineACC-LLM and ShapeACC-LLM) based on Qwen-VL (Qwen-VL)  [^2] Due to limited access, we do not employ popular MLLMs such as GPT-4o or Claude4-Sonnet.. Tab. 10 serves as an example. It can be observed that MLLMs are capable of specifically scoring the Line Straightness and Shape Congruence of the input images in accordance with the instructions of the prompts. Due to the characteristics of LLM, these type of metrics are sensitive to prompts, and there is also a possibility of hallucinations. However, it can still objectively reflect the quality of the rectification results to a certain extent.
@@
For the data utilized in the paper, our framework can also be applied on real-world test images. Considering that the portrait correction dataset (tan2021practical) is real-world and the remaining tasks are based on synthetic data, we focus on testing (1) rectified wide-angle image rectangling, (2) stitched image rectangling, and (3) rotation correction under real-world scenes as depicted in Fig. 16.
To evaluate the generalization ability for rectified wide-angle image rectangling, we collect rectified wide-angle image results by recent rectification methods (yang2021progressively; liao2021adeep), which includes various real-world wide-angle lenses. Fig. 16.1 shows that our method can also be well adapted to these real-world images. The original wide-angle images have obvious distortion while our results have reduced distortion and keep the rectangular borders.
For stitched image rectangling, we collect typical image stitching dataset (zaragoza2013projective; zara2014) and apply different image stitching methods  [^3] https://www.microsoft.com/en-us/research/product/computational-photography-applications/ to obtain panorama with seamless high-resolution. After getting the stitched images, we then perform rectangling as seen in Fig. 16.2. We have acheived the best rectangling visual effects with fewer distortions.
For rotation correction, we directly perform our method on the images from MS-COCO (lin2014microsoft) and the corrected results are shown in Fig. 16.3. These images have different resolutions and light conditions. However, our method can still perform well.
@@
Recent work MOWA (liao2024mowa) seems similar to this work. However, our motivations and structures differ in several ways: (1) Different tasks: Our model can perform all the tasks their model addresses, but their model does not consider rotation correction. (2) Unified framework: We provide a mathematical formulation of the general distortion model, whereas they do not offer any definition. (3) Network Architecture: Our network is aimed at our proposed general distortion model and based on up-to-date Mamba blocks.

### v2

@@ Appendix A Supplementary Materials for UniRect
In this section, we discuss the various tasks under our proposed general distortion model. The general distortion model incorporates the Kannala-Brandt model (Kannala and Brandt 2006), typically used to approximate wide-angle lenses, and the Brown-Conrady model (Weng et al. 1992), employed for radial distortion modeling. Consequently, we can employ this proposed distortion model to simulate real image distortion in real-world smartphone camera systems. As shown in Eq. (3), This model can be simplified as
@@
Optical flow is a concept related to the analysis of the movement of objects in a sequence of images. We utilize this tool to explore the distortion process for all pixels and formulate each distortion for every task. This is a qualitative tool. The generation of optical flow via different methods is presented in Fig. 11. Taking the distortion $\mathcal{T}_{2}^{-1}$ as an example, we present the results of various optical flow estimation methods, such as FlowNet (Dosovitskiy et al. 2015), FlowNet2 (Ilg et al. 2017), PWC-Net (Sun et al. 2018), and RAFT (Teed and Deng 2020). It can be observed from the figure that the outcomes of different optical flow estimation methods are similar without altering the mathematical form of distortion derived in this paper.
@@
Dataset Construction. For portrait correction $\mathcal{T}_{1}$, the data was collected using five different ultra wide-angle smartphone cameras. Over ten people were photographed in several scenes (Tan et al. 2021). The wide - angle portraits were manually corrected into distortion-free images. For the rectified wide-angle image rectangling dataset (Liao et al. 2023), wide-angle images were first synthesized, and the ground-truth was obtained manually. In the stitched image rectangling dataset (Nie et al. 2022) utilized for $\mathcal{T}_{3}$, the majority of the inputs are real stitching results, while the ground - truth is synthesized manually. For rotation correction $\mathcal{T}_{4}$ (Nie et al. 2023), the images are sourced from ImageNet (Deng et al. 2009) and processed manually.
@@
Zhu et al. (Zhu et al. 2022) $\mathcal{T}_{1}$ 8.79M - -
RecRecNet (Liao et al. 2023) $\mathcal{T}_{2}$ 62.70M - -
Nie et al. (Nie et al. 2022) $\mathcal{T}_{3}$ 52.14M - 20
RecDiffusion (Zhou et al. 2024) $\mathcal{T}_{3}$ 800M >1T -
DRC (Nie et al. 2023) $\mathcal{T}_{4}$ - - 5
@@
Here, we introduce the details of two metrics (i.e. LineACC-LLM and ShapeACC-LLM) based on Qwen-VL (Bai et al. 2023)  [^2] Due to limited access, we do not employ popular MLLMs such as GPT-4o or Claude4-Sonnet.. Tab. 10 serves as an example. It can be observed that MLLMs are capable of specifically scoring the Line Straightness and Shape Congruence of the input images in accordance with the instructions of the prompts. Due to the characteristics of LLM, these type of metrics are sensitive to prompts, and there is also a possibility of hallucinations. However, it can still objectively reflect the quality of the rectification results to a certain extent.
@@
For the data utilized in the paper, our framework can also be applied on real-world test images. Considering that the portrait correction dataset (Tan et al. 2021) is real-world and the remaining tasks are based on synthetic data, we focus on testing (1) rectified wide-angle image rectangling, (2) stitched image rectangling, and (3) rotation correction under real-world scenes as depicted in Fig. 16.
To evaluate the generalization ability for rectified wide-angle image rectangling, we collect rectified wide-angle image results by recent rectification methods (Yang et al. 2021; Liao et al. 2021b), which includes various real-world wide-angle lenses. Fig. 16.1 shows that our method can also be well adapted to these real-world images. The original wide-angle images have obvious distortion while our results have reduced distortion and keep the rectangular borders.
For stitched image rectangling, we collect typical image stitching dataset (Zaragoza et al. 2013, 2014) and apply different image stitching methods  [^3] https://www.microsoft.com/en-us/research/product/computational-photography-applications/ to obtain panorama with seamless high-resolution. After getting the stitched images, we then perform rectangling as seen in Fig. 16.2. We have acheived the best rectangling visual effects with fewer distortions.
For rotation correction, we directly perform our method on the images from MS-COCO (Lin et al. 2014) and the corrected results are shown in Fig. 16.3. These images have different resolutions and light conditions. However, our method can still perform well.
@@
Recent work MOWA (Liao et al. 2025) seems similar to this work. However, our motivations and structures differ in several ways: (1) Different tasks: Our model can perform all the tasks their model addresses, but their model does not consider rotation correction. (2) Unified framework: We provide a mathematical formulation of the general distortion model, whereas they do not offer any definition. (3) Network Architecture: Our network is aimed at our proposed general distortion model and based on up-to-date Mamba blocks.